The width of a rectangle is 5 less than twice its length. If the area of the rectangle is 194 cm^2, what is the length of the diagonal?
Answer by jorel1380(3719)

Let c be the length of the rectangle. Then its' width would be 2c-5. So:
c(2c-5)=194
2c²-5c-194=0
Using the quadratic formula, we get a positive value for c of 11.1779 for the length. Then the width would be 17.3557. Let d be the diagonal. Then:
(11.1779)²+(17.3557)²=d²
Solve for d
☺☺☺☺